OpenStudy (anonymous):

Im trying to finish some algebra so im not behind please help, Will give medals and fans to those who help....

OpenStudy (anonymous):

5.Create your own equation written in standard form. (a)Determine any point that is a solution and will be on this line. Justify your answer algebraically. (b)Explain how to find the x- and y- intercept(s) using complete sentences. Write the intercepts as ordered pairs. (c)Rewrite the equation using function notation. Graph the function and show the intercepts. 6.Compare the graph from question 2 and the equation from question 5. How are the graphs similar and different? Determine which has the greater y-value when x is 160 and explain how you determined it.

OpenStudy (anonymous):

Write an equation for a line such as... y = 2x + 4 This pick a point, like (1, 4) Plug it into equation (4) = 2(1) + 4 4 = 2 + 4 4 = 6 Not correct, so (1,4) isn't a point on the line

OpenStudy (anonymous):

Thank you very much :D @MarziaJacksonCute